OpenStudy (anonymous):

which are true? I know 12/8 is one but what is the other?

OpenStudy (anonymous):

@dan815 ?

OpenStudy (studygurl14):

What do you know about trigonometric ratios and the Pythagorean theorem?

OpenStudy (anonymous):

not alot I get confused easily @StudyGurl14

OpenStudy (studygurl14):

\(\large tan(\theta)=\Large\frac{side~opposite~angle}{side~adjacent~angle}=\frac{15}{8}\) 

OpenStudy (studygurl14):

To find the values of \(\large sin(\theta)\) and \(\large cos(\theta)\), you need to find the length of the hypotenuse.

OpenStudy (anonymous):

i'mconfused @Studylight

OpenStudy (studygurl14):

You know the lengths of two sides of the triangle, but you need to find the length of the hypotenuse. Use the Pythagorean Theorem \(\large a^2+b^2=c^2\rightarrow (8)^2+(15)^2=c^2\rightarrow c=\sqrt{64+225}\rightarrow c=\sqrt{289}=17\)

OpenStudy (studygurl14):

(cut off) \(\large\rightarrow c=\sqrt{289}=17=\tt the~hypotenuse\)

OpenStudy (studygurl14):

So... \(\large sin(\theta)=\frac{15}{17}\) and \(\large cos(\theta)=\frac{8}{17}\)

OpenStudy (studygurl14):

And.... \(\large sec(\theta)=\frac{17}{15}\) and... \(\large csc(\theta)=\frac{17}{8}\)

OpenStudy (studygurl14):

And... \(\large cot(\theta)=\frac{8}{15}\)

OpenStudy (studygurl14):

@KierseyClemons